Circular No. 8204

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2003 No. 2 N. J. Brown, Quinns Rocks, W. Australia, reports his discovery of a

possible nova (mag 9.2) on T-Max 400 film exposed with a 135-mm f/2 camera lens on Sept. 17.52 UT, providing a position of α = 18^{h}10^{m}13^{s}, δ = −27^{o}45′39′′ (equinox 2000.0). On Sept. 18.43, Brown estimated visual mag 8.8. H. Yamaoka, Kyushu University, reports that the new star appears at V ∼ 10.7 on an ASAS-3 CCD frame taken on Sept. 16.07 (though its image is merged with that of two neighboring stars). J. D. West, Mulvane, KS, reports the following precise position from four CCD V images (which yielded V = 8.98 ± 0.03, R_{c} = 8.49 ± 0.05, and [Hα] = 1.80 ± 0.03 for the nova on Sept. 19.085): α = 18^{h}10^{m}10^{s}.42, δ = −27^{o}45′35′′.2 (± 0′′.17). West adds that nothing is located at this position on a SERC V plate taken in 1987. W. Liller, Viña del Mar, Chile, reports that a grating spectrum (dispersion 0.78 nm/pixel) shows a prominent Hα emission (FWHM = 550 km/s) with a peak ≈ 2.3 times brighter than the surrounding continuum; a narrow P-Cyg absorption appears to be present and shifted 590 km/s towards the blue. CCD magnitudes from Liller (through thin cirrus): Sept. 19.104, V = 8.91 ± 0.05; 19.106, B = 9.79 ± 0.15. Nothing is visible to mag 11.5 on a red film taken by Liller on Sept. 13.12.

G. Ruch, E. Polomski, and C. E. Woodward, University of Minnesota; R. M. Wagner, LBT Observatory; and S. Starrfield, Arizona State University, report that optical spectra (range 370–640 nm, resolution 0.6 nm) were obtained with the Steward Observatory Bok 2.3-m telescope on Kitt Peak on Sept. 19.10 UT, showing that the spectrum of Brown’s object exhibits emission lines of Hβ (equivalent width 0.6 nm), Hγ, Hδ and numerous Fe II lines with P-Cyg profiles superposed on a flat continuum. The FWHM of Hβ emission is 612 km/s, and the terminal velocity of the absorption component is 800 km/s. The equivalent width of the diffuse interstellar line at 578 nm is 0.05 nm, corresponding to E(B − V) ∼ 0.1 mag. The spectra confirm that this object is an Fe-II-class nova in its early decline stage.

SUPERNOVA 2003ia IN NGC 6109 Further to IAUC 8203, T. Matheson et al. report that a spectrum of

SN 2003ia (cf. IAUC 8201), obtained by Calkins on Sept. 19.18 UT, shows it to be a type-Ia supernova, ∼ 2 weeks past maximum.

2003 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green